OpenStudy (anonymous):

Solve the equation: 4x2 = 9

OpenStudy (anonymous):

Factor: 9x2 - 24x + 16

OpenStudy (anonymous):

9x2 - 24x + 16 = x=4/3

OpenStudy (anonymous):

x=4/3